# AC-Lite : A Lightweight Image Captioning Model for Low-Resource Assamese Language

### v1

@@ 4.1 Hyper Parameters
The visual feature encoder takes RGB images of size $3\times 224\times 224$ as input and produces a feature map of size $1024\times 14\times 14$. These features are extracted from the final convolutional layers of the CNN. Later, adaptive pooling transforms this feature map into $n_{a}=14\times 14=196$ feature vectors of $d_{a}=1024$ dimension. The GRU and attention GRU hidden layer sizes are set to $\mathbf{h}_{t}^{G}=\mathbf{h}_{t}^{A}=512$. The attention embedding dimension is set to $d_{e}=512$. The model is trained by using Adam optimizer with a learning rate of $5\times 10^{-4}$. The model is trained for 30 epochs with a batch size of 50 image-caption pairs.
@@ 4.3 Baseline Models
Baseline-4: The Baseline-4 employs the Object Relation Transform (ORT) model by Herdade et al. [6]. This model uses the geometric position of objects in the transformer encoder instead of positional encoding. These geometric features adjust attention weights and improve the ability of the model to capture spatial relationships within images for more accurate caption generation.
Baseline-5: The Baseline-5 employs the Attention on Attention Network (AoANet) by Huang et al. [5]. This model uses the Attention on Attention (AoA) module, which consists of a multi-head self-attention followed by an additional attention layer. The AoA module is used for visual feature refinement and with the LSTM decoder for caption generation.
Baseline-6: The Baseline-6 employs a Bi-LSTM network with Bilinear Attention [13]. The Bi-LSTM captures sequential dependencies in both forward and backward directions which is important for capturing the linguistic properties of Assamese language. Additionally, the bilinear attention applies second-order interactions to improve model performance.
Table 2: Performance of the proposed AC-Lite model with baseline and State-of-the-art Assamese captioning model trained and tested on the COCO-AC dataset. MFLOPs imply Million FLOPs, and parameters are measured in Millions.
Models BLEU-1 BLEU-2 BLEU-3 BLEU-4 CIDEr FLOPs (M) Param (M)
Baseline-1 59.9 41.5 29.1 20.6 61.5 2202.21 26.11
Baseline-2 62.1 45.1 31.8 23.2 68.3 16192 158.46
Baseline-3 67.1 49.6 36.5 27 80.8 119452.92 122.6
Baseline-4 66.3 48.8 35.7 26.2 79.7 135797.2 119.35
Baseline-5 67.1 50.1 37 27.1 82.4 119230.24 119.25
Baseline-6 68.1 50.5 37.2 27.5 81.7 121580.54 181.7
AC-Lite + XE 63.6 45.8 33.1 24.1 72 1097.57 25.65
AC-Lite + RL 68.3 50.2 36.2 26 82.3 1097.57 25.65
Table 3: Performance of the proposed AC-Lite model with baseline and State-of-the-art Assamese captioning model trained and tested on the Flickr30K-AC dataset. MFLOPs imply Million FLOPs, and parameters are measured in Millions.
Models BLEU-1 BLEU-2 BLEU-3 BLEU-4 CIDEr FLOPs (M) Param (M)
Baseline-1 51.1 32.5 21 13.3 31.9 2202.21 26.11
Baseline-2 53.3 35.6 24.1 16.5 39.7 16192 158.46
Baseline-3 60.4 42.3 29.5 20.1 46.5 119452.92 122.6
Baseline-4 60.1 41.8 28.7 20 47.2 135797.2 119.35
Baseline-5 61 43 30.1 21.2 50.4 119230.24 119.25
Baseline-6 61.5 43.4 30.4 21.1 48.3 121580.54 181.7
AC-Lite + XE 54.9 37 25.1 16.6 40 1097.57 25.65
AC-Lite + RL 58.5 40.2 27.9 18.4 46.1 1097.57 25.65
@@ 5.1 Quantitative Results
The quantitative results are obtained by using beam sizes of 6 and 3 for COCO-AC and Flickr30K-AC, respectively. The Table 2 and Table 3 compare the performance of the proposed AC-Lite model with the existing Assamese image captioning model and baseline models. The AC-Lite model trained with cross-entropy loss (AC-Lite + XE) achieves respective BLEU-1, BLEU-4, and CIDEr scores of 63.6, 24.1, and 72 on COCO-AC, and 54.9, 16.6, and 40.0 on Flickr30K-AC, with only 1097.57 MFLOPs and 25.65M parameters. Compared to Baseline-5 and Baseline-6, the performance of the AC-Lite model is slightly lower. However, the computation requirement is 110$\times$ lower than state-of-art Assamese captioning models. Moreover, when the proposed model is trained with reinforcement learning (AC-Lite + RL), it outperforms all other models, achieving 68.3 BLEU-1, 26 BLEU-4, and 82.3 CIDEr on COCO-AC. In the case of Flickr30K-AC, the model trained with reinforcement learning (AC-Lite + RL) gives competitive results achieving 58.5 BLEU-1, 18.4 BLEU-4, and 46.1 CIDEr. This impressive performance with less computation indicates that the AC-Lite model is suitable for deployment in resource-constrained environments.
Table 4: Performance of the proposed AC-Lite model with different pretrained Convolutional Neural Networks used as visual feature encoder. All models are trained and tested on the COCO-AC dataset. MFLOPs imply Million FLOPs, and parameters are measured in Millions. Note– B-1: BLEU-1, B-4: BLEU-4 and C: CIDEr.
@@
Visual Feature Encoder BLEU-1 BLEU-2 BLEU-3 BLEU-4 CIDEr FLOPs (M) Param (M) ShuffleNetV2x0.5 57.7 39.6 27.6 19.3 59.1 837.575 23.55 MobileNetV3_Small 56.6 38.5 26.1 17.8 55.6 857.575 24.65 MobileNetV1X0.25 56.5 38.6 26.8 18.8 55.7 886.715 22.62 ShuffleNetV2x1.0 61.1 43.3 30.8 22 66.2 937.575 24.45 MobileNetV3_Large 59.5 41.6 28.7 19.9 61.6 1017.575 27.65 ShuffleNetV2x1.5 63.6 45.8 33.1 24.1 72 1097.575 25.65 EfficientNetB0 59.6 42 29.1 20.1 61.6 1187.575 27.45 ShuffleNetV2x2.0 62.8 45.2 32.5 23.4 70.5 1377.575 29.55 EfficientNetB1 59.6 41.9 29 20.2 63.1 1487.575 29.95 ResNet101 66.4 48.8 35.7 26.1 78.5 8597.575 66.65
@@ 5.2 Ablation Analysis
The Table 4 presents an ablation study evaluating different CNN architectures as visual feature encoders for the proposed model. The experiments aim to identify the encoder that provides the best performance while maintaining low computational requirements. Among lightweight CNN architectures, ShuffleNetv2x1.5 achieves the best performance, with BLEU-1, BLEU-4, and CIDEr scores of 63.6, 24.1, and 72.0, respectively, at computation costs of 1097.57 MFLOPs and 25.65M parameters. ShuffleNetv2x1.0 and ShuffleNetv2x2.0 also perform well, with CIDEr scores of 66.2 and 70.5, respectively, demonstrating the efficiency of the ShuffleNetv2 family. EfficientNetB0 and EfficientNetB1 yield moderate results but require higher FLOPs and parameters compared to ShuffleNet variants. ResNet101 achieves the highest performance, with a CIDEr score of 78.5, but its computational cost is significantly higher, requiring 8597.57 Million FLOPs and 66.65 Million parameters. This analysis indicates that ShuffleNetv2x1.5 provides an effective trade-off between performance and computational efficiency, making it a suitable choice of encoder for lightweight image captioning models.
@@ 5.3 Qualitative Analysis
The Figure 3 shows qualitative examples of captions generated by the proposed AC-Lite model for the test set of the COCO-AC dataset. The Figure 3 compares the quality of the generated captions of AC-Lite with those reported in BiLSTM [13]. For example, as illustrated in Figure 3(a) and Figure 3(b), the proposed model generates similar or better captions compared to Baseline-6 [13]. These generated captions can identify the color of the shirt as [Uncaptioned image] (Orange) as well as activity [Uncaptioned image] (Snowboarding) and [Uncaptioned image] (Playing Tennis). Additionally, in Figure 3(a), the proposed model can identify the environment as [Uncaptioned image] (Mountain) more accurately. The same can be observed with Figure 3(b) and Figure 3(f) where the model accurately identifies the surrounding environment [Uncaptioned image] (Sidewalk) and [Uncaptioned image] (City) respectively. This is attributed to the fact that the CNN models provide the whole image as input to the language decoder instead of a few selected image regions. However, inputting all image regions has drawbacks, such as the proposed model failing in object counting and identification. For instance, in Figure 3(c) counts one person. Similarly, the proposed model fails to identify the “horse” and “car” present in the Figure 3(d) and Figure 3(f), respectively.
Figure 3: Qualitative example produced by the proposed AC-Lite on COCO-AC test set. Here, AC-Lite – caption generated by the proposed model, BiLSTM – caption reported in  [13], gloss – gloss annotation.
@@ 6 Conclusion
This paper introduces AC-Lite, a lightweight image captioning model designed for the Assamese language. The AC-Lite model is designed to meet the need for computationally efficient solutions that can operate on resource-constrained devices. AC-Lite employs lightweight ShuffleNetv2x1.5 for visual feature extraction and Gated Recurrent Units (GRUs) in the decoder. This incorporation significantly reduces computational demands and model parameters. Additionally, the integration of bilinear attention further enhances captioning performance. AC-Lite is designed to function independently on device hardware, thereby eliminating dependency on remote servers. This makes it suitable for real-world applications like assistive technologies and smart education technology for Assamese speakers, particularly in rural areas. Future efforts will focus on improving semantic understanding and further optimizing the model for enhanced performance in diverse real-world scenarios.

### v2

@@ 4.1 Hyper Parameters
The visual feature encoder takes RGB images of size $3\times 640\times 480$ as input and produces a feature map of size $1024\times 14\times 14$. These features are extracted from the final convolutional layers of the CNN. Later, adaptive pooling transforms this feature map into $n_{a}=14\times 14=196$ feature vectors of $d_{a}=1024$ dimension. The GRU and attention GRU ($\mathbf{GRU_{A}}$) hidden layer sizes are set to $\mathbf{h}_{t}^{G}=\mathbf{h}_{t}^{A}=512$. The attention embedding dimension is set to $d_{e}=512$. The model is trained by using Adam optimizer with a learning rate of $5\times 10^{-4}$. The model is trained for 30 epochs with a batch size of 50 image-caption pairs.
@@ 4.3 Baseline Models
AC-Lite is compared against the following nine baseline models.
@@
Baseline-4: The Baseline-4 employs the Attention on Attention Network (AoAN-et) by Huang et al. [9]. This model uses the Attention on Attention (AoA) module, which consists of a multi-head self-attention followed by an additional attention layer. The AoA module is used for visual feature refinement and with the LSTM decoder for caption generation.
Baseline-5: The Baseline-5 employs the Object Relation Transform (ORT) model by Herdade et al. [10]. This model uses the geometric position of objects in the transformer encoder instead of positional encoding. These geometric features adjust attention weights and improve the ability of the model to capture spatial relationships within images for more accurate caption generation.
Baseline-6: The Baseline-6 uses the Mesh-Memory Transformer by Cornia et al. [19], which connects encoder and decoder sub-layers in a mesh structure to capture multi-level visual relationships. It serves to evaluate whether integrating semantic information offers improvements over memory-based connections.
Baseline-7: The Baseline-7 employs the standard transformer model introduced by Vaswani et al. [35], selected for its foundational role in sequence generation. It uses an encoder to refine visual features, which are then decoded into captions. Six layers of encoders and decoders are stacked to facilitate fair comparison with models incorporating semantic and spatial enhancements.
Baseline-8: The Baseline-8 employs a Bi-LSTM network with Bilinear Attention [17] for Assamese image caption generation. The Bi-LSTM captures sequential dependencies in both forward and backward directions which is important for capturing the linguistic properties of Assamese language. Additionally, the bilinear attention applies second-order interactions to improve model performance.
Baseline-9: The Baseline-9 integrates semantic attributes with visual features [5] using VSSA module for Assamese image caption generation. The VSSA module uses separate self-attention submodule to process semantic attributes with visual features. Additionally, a Bi-GRU is employed to capture both forward and backward contextual information, which is particularly important for modeling the linguistic characteristics of the Assamese language.
@@ 5.1 Quantitative Results
The quantitative results are obtained by using beam sizes of 6 and 3 for COCO-AC and Flickr30K-AC, respectively. The Table 2 and Table 3 compare the performance of the proposed AC-Lite model with baseline models. Note that the computation-complexity and number of parameters are different for the same model in Table 2 and Table 3 because of the different vocabulary sizes for COCO-AC and Flickr30K-AC.
Table 2: Performance of AC-Lite model compared against nine baselines. All models are trained and tested on the COCO-AC dataset. AC-Lite+XE and AC-Lite+RL respectively denote cross-entropy loss minimization and reinforcement learning based model training. Performances are reported in terms of BLEU-n, CIDEr scores, computations (GFLOPs) and parameters in millions (M).
Models BLEU-1 BLEU-2 BLEU-3 BLEU-4 CIDEr GFLOPs Params (M) Baseline-1 59.9 41.5 29.1 20.6 61.5 48.121 59.41 Baseline-2 62.1 45.1 31.8 23.2 68.3 48.392 60.46 Baseline-3 67.1 49.6 36.5 27 80.8 118.336 122.61 Baseline-4 67.3 50.1 37 27.6 81.5 122.839 156.21 Baseline-5 66.3 48.8 35.7 26.2 79.7 120.105 137.47 Baseline-6 66.9 49.5 36.6 27.4 80.4 119.398 112.67 Baseline-7 67.4 49.6 36.3 26.8 81.1 120.227 122.06 Baseline-8 68.1 50.5 37.2 27.5 81.7 119.578 173.72 Baseline-9 68.6 51 38.1 28.6 84.5 120.01 178.04 AC-Lite + XE 63.6 45.8 33.1 24.1 72 2.45 22.87 AC-Lite + RL 68.3 50.2 36.2 26 82.3 2.45 22.87
Table 3: Performance of AC-Lite model compared against nine baselines. All models are trained and tested on the Flickr30K-AC dataset. AC-Lite+XE and AC-Lite+RL respectively denote cross-entropy loss minimization and reinforcement learning based model training. Performances are reported in terms of BLEU-n, CIDEr scores, computations (GFLOPs) and parameters in millions (M).
Models BLEU-1 BLEU-2 BLEU-3 BLEU-4 CIDEr GFLOPs Params (M) Baseline-1 51.1 32.5 21 13.3 31.9 48.081 54.92 Baseline-2 53.3 35.6 24.1 16.5 39.7 48.352 55.98 Baseline-3 60.4 42.3 29.5 20.1 46.5 118.258 113.85 Baseline-4 61 42.7 30.1 21.2 48.6 122.825 147.24 Baseline-5 60.1 41.8 28.7 20 47.2 120.013 115.47 Baseline-6 59.5 42.1 29.4 20.1 47.4 119.332 100.36 Baseline-7 60.6 43 30.3 20.9 48 120.186 117.58 Baseline-8 61.5 43.4 30.4 21.1 48.3 119.42 164.2 Baseline-9 62.7 45.2 32.3 23 51.4 119.937 167.77 AC-Lite + XE 54.9 37 25.1 16.6 40 2.41 18.38 AC-Lite + RL 58.5 40.2 27.9 18.4 46.1 2.41 18.38
@@
As illustrated in Table 2, the AC-Lite model trained with cross-entropy loss (AC-Lite + XE) achieves respective BLEU-1, BLEU-4, and CIDEr scores of 63.6, 24.1, and 72 on COCO-AC with only 2.45 GFLOPs and 22.87M parameters. For Flickr30K-AC, the AC-Lite model achieves 54.9 BLEU-1, 16.6 BLEU-4, and 40.0 CIDEr score with only 2.41 GFLOPs and 18.38M parameters. Compared to specifically designed Assamese image captioning models Baseline-8 and Baseline-9, the performance of the AC-Lite model is slightly lower. However, the computation requirement is approximately 48$\times$ lower than these Assamese captioning models. Moreover, when the proposed AC-Lite is trained with reinforcement learning (AC-Lite + RL) achieving 68.3 BLEU-1, 26 BLEU-4, and 82.3 CIDEr on COCO-AC. In the case of Flickr30K-AC, the model trained with reinforcement learning (AC-Lite + RL) gives competitive results achieving 58.5 BLEU-1, 18.4 BLEU-4, and 46.1 CIDEr. This highlights that AC-Lite achieves a favorable trade-off between performance and efficiency, making it suitable for deployment in resource-constrained applications.
Table 4: Performance of the proposed model with different pretrained CNN backbone based visual feature encoders. GRUs are used for both attention and decoder units. All models are trained and tested on the COCO-AC dataset by cross-entropy loss minimization. Performances are reported in terms of BLEU-n, CIDEr scores, computations (GFLOPs) and parameters in millions (M).
@@
Visual Feature Encoder BLEU-1 BLEU-2 BLEU-3 BLEU-4 CIDEr GFLOPs Params (M) ShuffleNetv2x0.5 57.7 39.6 27.6 19.3 59.1 0.865 20.730 MobileNetV3_Small 56.6 38.5 26.1 17.8 55.6 0.964 21.32 ShuffleNetv2x1.0 61.1 43.3 30.8 22 66.2 1.516 21.64 MobileNetV3_Large 59.5 41.6 28.7 19.9 61.6 1.972 23.36 ShuffleNetv2x1.5 63.6 45.8 33.1 24.1 72 2.45 22.87 MobileNetV2 56.5 38.6 26.8 18.8 55.7 2.527 22.610 EfficientNetB0 59.6 42 29.1 20.1 61.6 3.059 24.4 ShuffleNetv2x2.0 62.8 45.2 32.5 23.4 70.5 4.216 25.73 EfficientNetB1 59.6 41.9 29 20.2 63.1 4.221 26.9 ResNet18 63.4 45.7 32.9 23.8 72.2 11.745 31.57 ResNet50 65.7 48.1 35 25.6 76.8 25.771 43.9 ResNet101 66.4 48.8 35.7 26.1 78.5 48.562 62.89
@@ 5.2 Ablation Analysis
The Table 4 presents an ablation study evaluating different CNN architectures as visual feature encoders for the proposed model. The experiments aim to identify the encoder that provides the best performance while maintaining low computational requirements. Among lightweight CNN architectures, ShuffleNetv2x1.5 achieves the best performance, with BLEU-1, BLEU-4, and CIDEr scores of 63.6, 24.1, and 72.0, respectively, at total computation costs of 2.45 GFLOPs and 22.87M parameters. ShuffleNetv2x1.0 and ShuffleNetv2x2.0 also perform well, with CIDEr scores of 66.2 and 70.5, respectively, demonstrating the efficiency of the ShuffleNetv2 family. EfficientNetB0 and EfficientNetB1 yield moderate results but require higher FLOPs and parameters compared to ShuffleNet variants. ResNet101 achieves the highest performance, with a CIDEr score of 78.5, but its computational cost is significantly higher, requiring 48.581 GFLOPs and 63.94 million parameters. This analysis indicates that ShuffleNetv2x1.5 provides an effective trade-off between performance and computational efficiency, making it a viable encoder option for lightweight image captioning models.
Table 5: Performance of the proposed model with different RNNs for attention and language decoding. ShuffleNetV2x1.5 is used as the visual feature encoder. The model is trained and tested on the COCO-AC dataset by cross-entropy loss minimization. Performances are reported in terms of BLEU-n, CIDEr scores, computations (GFLOPs) and parameters in millions (M).
Attention RNN Language RNN BLEU-1 BLEU-4 CIDEr GFLOPs Params (M) LSTM LSTM 62.8 23.9 71 2.483 24.71 LSTM GRU 63 23.5 69.8 2.469 23.92 GRU LSTM 62.7 23.4 70 2.464 23.65 GRU GRU 63.6 24.1 72 2.45 22.87
Figure 3: Qualitative example produced by the proposed AC-Lite on COCO-AC test set. Here, AC-Lite – caption generated by the proposed model, gloss – gloss annotation.
Another ablation study was conducted to assess the impact of different types of RNNs used in the attention and language modules of the proposed model. For this ablation study all experiments were performed using ShuffleNetV2x1.5 as the visual feature encoder and evaluated on the COCO-AC dataset. As shown in Table 5, the combination of GRU for both the attention and language modules yields the best performance, achieving the highest BLEU-1 as 63.6, BLEU-4 as 24.1, and CIDEr score 72, while also maintaining the lowest computational cost at 2.45 GFLOPs and 22.87 million parameters. In comparison, replacing either GRU with LSTM leads to drops in performance and increased computational load. Notably, the LSTM-LSTM configuration, while achieving a slightly lower CIDEr score of 71, is the most resource-intensive with 2.483 GFLOPs and 24.71M parameters. These results indicate that the GRU-GRU combination improves efficiency and enhances the overall captioning quality, making them the preferred choice for lightweight and effective image captioning models.
@@ 5.3 Qualitative Analysis
The Figure 3 shows qualitative examples of captions generated by the proposed AC-Lite model for the test set of the COCO-AC dataset. For example, as illustrated in Figure 3(a) and Figure 3(b), the AC-Lite model can identify the color of the shirt as [Uncaptioned image] (Orange) as well as activity [Uncaptioned image] (Snowboarding) and [Uncaptioned image] (Playing Tennis). Additionally, in Figure 3(a), the proposed model can identify the environment as [Uncaptioned image] (Mountain) more accurately. The same can be observed with Figure 3(b) and Figure 3(f) where the model accurately identifies the surrounding environment [Uncaptioned image] (Sidewalk) and [Uncaptioned image] (City) respectively. This is attributed to the fact that the CNN models provide the whole image as input to the language decoder instead of a few selected image regions. However, inputting all image regions has drawbacks, such as the proposed model failing in object counting and identification. For instance, in Figure 3(d), the proposed model counts one person. Similarly, the proposed model fails to identify the “horse” and “car” present in the Figure 3(e) and Figure 3(f), respectively.
@@ 6 Conclusion
This paper introduces AC-Lite, a lightweight image captioning model designed for the low-resource Assamese language. The AC-Lite model is designed to meet the need for computationally efficient solutions that can operate on resource-constrained devices. AC-Lite employs lightweight ShuffleNetv2x1.5 for visual feature extraction and Gated Recurrent Units (GRUs) in the decoder. This incorporation significantly reduces computational demands and model parameters. Additionally, the integration of bilinear attention further enhances captioning performance. AC-Lite is designed to function independently on device hardware, thereby eliminating dependency on remote servers. This makes it suitable for real-world applications like assistive technologies and smart education technology for Assamese speakers, particularly in remote areas having issues with continuous high-speed internet connectivity.
Future efforts will primarily focus on the following important directions. First, the proposal of novel lightweight image feature extractor and language decoder architectures along with efficient training strategies. Second, the development of application-specific image-caption datasets for assistive healthcare and education technologies in Indian contexts. Finally, system-level implementation of the lightweight image captioning model on resource-constrained edge devices for application specific deployment.